40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7516     Entry time: Wed Oct 10 02:20:34 2012
Author: rana 
Type: Update 
Category: SUS 
Subject: Optical Lever QPD mods 

 Since we upgraded the CDS system, I guess our ADC ranges have gone up but we never did anything to the OLs to match the ADC ranges. From Liz's daily summary page of the OL, I see this:

C1-CORE_OPTICS_SUM_DATA_ALL_TIME_0-1033801216-86400.png

So we need a factor of 5-10 increase in the electronics gain (why isn't the BS SUM on there?). This might be accomplished in the head, but for the ones with whitening boards, might be better to do there.

(** add to Jamie's list of long term tasks **)

ELOG V3.1.3-